# Quantum loops in the 1T-transition metal dichalcogenides

### v1

@@ Abstract
Introduction—  Loops and loop-superpositions appear in many exciting contexts. For example, Kitaev’s toric code has a spin-liquid ground state, arising from a quantum superposition of all possible ways of drawing loops on the square latticeKitaev (2003); Trebst et al. (2007). A second example is spin-ice where each allowed ice state can be viewed as a loop-covering of the diamond latticeJaubert et al. (2011); Gingras and McClarty (2014). Quantum loop models (QLMs) offer a paradigm to study such systemsRan et al. (2023, 2024). They describe dynamics within the set of all loop-coverings of the underlying lattice. Each loop covering is assigned a potential energy arising from proximity between loop segments. Dynamics arises from tunnelling processes where proximate loops are cut and crossed. Here, we motivate a QLM on physical grounds with a view to describe bonding in a class of transition metal dichalcogenides (TMDs).
The building block of TMDs is the X-M-X trilayer, where M and X are transition metal and chalcogenide atomsManzeli et al. (2017). Each layer forms a triangular lattice. When these three layers are ABC-stacked, the resulting structure is denoted as 1T and is prone to distortions. A prominent example is the 1T structure which exhibits superconductivityQi et al. (2016); Guo et al. (2017); Fang et al. (2018) and topological bandsQian et al. (2014); Shi et al. (2019). We build a model for bonding in 1T structures with a view to describing distortions.
@@
In the 1T crystal structure, the octahedral crystal field splits the $d$ orbitals into t2g and eg levelsPav (2023). In a d2 configuration, the two electrons occupy the lower-lying t2g orbitals. Crucially, in the undistorted 1T structure, the three t2g orbitals show strong directionality in overlaps between neighbouring atoms. To see this, we first note that $xy$, $yz$ and $zx$ orbitals are defined with respect to a local coordinate system at each transition metal atom. The $x$, $y$ and $z$ axes point towards chalcogen atoms that lie at the vertices of an octahedron as shown in Fig. 1(a). The six nearest neighbour vectors on the triangular lattice lie along ±$\pm(\hat{x}+\hat{y})$, ±$\pm(\hat{y}+\hat{z})$ and ±$\pm(\hat{z}+\hat{x})$. Along the ±$\pm(\hat{x}+\hat{y})$ directions, neighbouring atoms share the same $xy$ plane. Two such neighbours have a strong overlap between their dxy orbitals, and a zero overlap between any other pair of t2g orbitals. Similarly, the dzx and dyz orbitals have strong overlaps along ±$\pm(\hat{z}+\hat{x})$ and ±$\pm(\hat{y}+\hat{z})$ bonds respectively, as shown in Fig. 1(b-d).
@@
Quantum loop model: Hamiltonian— We now define a minimal Hamiltonian à la Rokhsar-KivelsonRokhsar and Kivelson (1988). We first define potential energy with two contributions:
@@
These energy contributions encode proximity of bonds. For example, electrons come closer when bonds form an acute angle rather than at an obtuse angle. This can impose an energy cost due to Coulomb repulsion, corresponding to a positive value of $V$. In a triangle loop, we have three bonds that are very close to one another. This could give rise to a local distortion with the three vertices coming closer to one another, with an associated energy cost $V^{\prime}$. We treat $V$ and $V^{\prime}$ as independent parameters that can each take positive or negative values.
@@
This Hamiltonian differs from previous studies that have adapted Rokhsar-Kivelson-like terms to a loop settingRan et al. (2023, 2024). There is a difference even at the level of the Hilbert space, due to the bending constraint at each site that arises from orbital character. This further constrains the kinetic energy term – a flipping process is only allowed if the initial and final configurations satisfy the bending constraint. The second key difference is in the form of the potential energy terms. Previous studies have defined potential energy in direct analogy with the original Rokhsar-Kivelson model, associating an energy cost with each flippable plaquette. Here, we have defined potential energy on physical grounds motivated by the geometry and physics of TMDs.
Numerical approach—  To solve for the ground state of Eq. 10, we first enumerate all loop coverings to construct the Hilbert space. This is a challenging task, with no known analytic solution. We use a numerical approach, working with finite clusters and periodic boundaries. We use a recursive branching algorithm to find all loop coverings that satisfy the bending constraint. The number of configurations grows rapidly with system size: 42 for 3x3, 250 964 for 5x4, 2 720 400 for 6x4. For concreteness, we present results from the 6x4 cluster below. We expect the same qualitative behaviour with larger sizes.
We construct the Hamiltonian as an 2 720 400 $\times$ 2 720 400 matrix, with flipping processes entering in off-diagonal terms. The resulting matrix is sparse, with $\sim 10^{-4}\%$ of the entries being non-zero. We use a Krylov-space based approach to find the lowest-lying eigenstates.
@@
Phase diagram—  We find the ground state as a function of three parameters, $t$, $V$ and $V^{\prime}$. Fig. 3(e,f) depicts two slices of the three-dimensional phase diagram. We find three phases:
@@
We identify the acute stripe phase with the 1T phaseYin et al. (2021); Tang et al. (2021); Dai et al. (2024) seen in many materials. Bond correlations in the QLM ground state show a clear stripe-like pattern. This pattern is very similar to the variation of bond lengths seen in 1T materials – see supplementsup . In the QLM ground state, a stronger bond is indicated by a higher likelihood of dimer occupancy. In materials, a stronger bond is one with a shorter bond length.
(ii) Obtuse stripe phase: This phase appears at positive values of $V$ where obtuse angles are favoured. Indeed, obtuse stripe configurations maximize the number of obtuse angles with all other configurations having fewer obtuse angles. They also allow for a high degree of dynamics with many flippable plaquettes. The ground state is a linear superposition with dominant weight from an obtuse stripe configuration.
(iii) Trimerized phase: We find a single-configuration-state that contains an arrangement of triangle-loops. On the 6$\times$4 lattice, we find a trimerized phase with alternating ‘up’ and ‘down’ triangles as shown in Fig. 3(b). On other sizes, we find configurations with all-up or all-down – see supplementary materialsup . Trimerized phases do not allow for flipping processes within our model. Dynamics may appear in a more general model with longer-range rearrangements.
A trimerized phase is known to appear in AVX2, where A=Li, Na and X=O,S,Si. Previous studies have described trimerization as arising from orbital orderingPen et al. (1997a, b); Ezhov, S. Yu. et al. (1998). Our QLM places the trimerized phase within the broader context of loop model phases.
Impurity textures—  Our QLM description is a real-space approach based on short-ranged valence bond formation. This can be contrasted with momentum-space-based studies, e.g., where distortions are viewed as exciton-like instabilitiesMonney et al. (2011); Kogar et al. (2017); Si et al. (2020) or as Peierls transitionsDuerloo et al. (2014); Keum et al. (2015); Li and Li (2016). To build a case for a real-space approach, we present a testable prediction. We consider a single impurity where a d1 transition metal substitutes at the site of the d2 metal atom. For example, we may have a Nb atom at a Mo site in 1T-MoS2. A d1 atom can only form a single valence bond. As a result, it cannot allow a loop to pass through. Rather, it serves as a loop-termination point.
@@
As a further test, we may consider a d3, rather than d1, impurity. This creates a T-junction rather than a loop termination point. This leads to precisely the same distortion patterns as d1, as shown in the supplementsup .
@@
Discussion—  We have proposed a QLM description for 1T TMDs, where distortions are driven by short-ranged valence bond formation. Loops emerge from d2 character and the directionality of t2g overlaps, with each loop containing a sequence of valence bonds. This picture is justified in a Mott insulator driven by strong intra-orbital Hubbard repulsion. If Hund’s coupling is weak, we may identify each valence bond as a covalent bond or a spin-singlet wavefunction. If Hund’s coupling is strong, two electrons on the same atom cannot form independent singlet bonds. Rather, electron spins will align to give rise to an effective spin-1 moment on the atom. Nevertheless, the QLM description can hold – if each loop is interpreted as an emergent Haldane chain. This picture has been evoked to described the trimerized phase of LiVO2Pen et al. (1997b). Based on the same idea, loop models have been proposed on the square lattice in the context of LaFeAsOBaskaran (2008) and more recently on the honeycomb latticeSavary (2021).
In MoX2 and WX2 (X=S,Se), the 1T structure is known to have a small electronic gapKeum et al. (2015); Chen et al. (2018). However, the 1T structure is metallicYin et al. (2021). Naively, this goes against our premise of localized electrons. If these distortion-free materials are described by a QLM, they may arise from a putative loop-liquid state. This state may host mobile charged excitations, reminiscent of holons and doublons in the the RVB theory of high-temperature superconductivity. This is an exciting direction for future studies, motivating the search for a quantum loop liquid in analogy with quantum spin liquids. Notably, a loop liquid phase has been seen in quantum Monte Carlo simulations of a related loop modelRan et al. (2024).
@@
Supplemental Material for “Quantum loops in the 1T-transition metal dichalcogenides”
A. Knowles,1, G. Baskaran2,3 and R. Ganesh1
@@
(Dated: September 5, 2024)
@@ III Trimerized phases
Figure S5: a) Dimer density in the acute stripe ground state of the QLM. The ground state is projected to only retain states where dimers are present at the two bonds at the bottom left. Bond colour and line width are proportional to the dimer probability. b) Bond lengths in the MoS2 using data from Ref. Qian et al., 2014. Darker, thicker bonds have shorter bond lengths while lighter, thinner bonds are longer.

### v2

# Quantum loops in the 1T transition metal dichalcogenides

@@ Abstract
Introduction—  Loops and loop-superpositions appear in many exciting contexts. For example, Kitaev’s toric code has a spin-liquid ground state, arising from a quantum superposition of all possible ways of drawing loops on the square lattice[1, 2]. A second example is spin-ice where each allowed ice state can be viewed as a loop-covering of the diamond lattice[3, 4]. Quantum loop models (QLMs) offer a paradigm to study such systems[5, 6]. They describe dynamics within the set of all loop-coverings of the underlying lattice. Each loop covering is assigned a potential energy arising from proximity between loop segments. Dynamics arises from tunnelling processes where proximate loops are cut and crossed. Here, we motivate a QLM on physical grounds with a view to describe bonding in a class of transition metal dichalcogenides (TMDs).
The building block of TMDs is the X-M-X trilayer, where M and X are transition metal and chalcogenide atoms[7]. Each layer forms a triangular lattice. When these three layers are ABC-stacked, the resulting structure is denoted as 1T and is prone to distortions. A prominent example is the 1T structure which exhibits superconductivity[8, 9, 10] and topological bands[11, 12]. We build a model for bonding in 1T structures with a view to describing distortions.
@@
In the 1T crystal structure, the octahedral crystal field splits the $d$ orbitals into t2g and eg levels[13]. In a d2 configuration, the two electrons occupy the lower-lying t2g orbitals. Crucially, in the undistorted 1T structure, the three t2g orbitals show strong directionality in overlaps between neighbouring atoms. To see this, we first note that $xy$, $yz$ and $zx$ orbitals are defined with respect to a local coordinate system at each transition metal atom. The $x$, $y$ and $z$ axes point towards chalcogen atoms that lie at the vertices of an octahedron as shown in Fig. 1(a). The six nearest neighbour vectors on the triangular lattice lie along ±$\pm(\hat{x}+\hat{y})$, ±$\pm(\hat{y}+\hat{z})$ and ±$\pm(\hat{z}+\hat{x})$. Along the ±$\pm(\hat{x}+\hat{y})$ directions, neighbouring atoms share the same $xy$ plane. Two such neighbours have a strong overlap between their dxy orbitals, and a zero overlap between any other pair of t2g orbitals. Similarly, the dzx and dyz orbitals have strong overlaps along ±$\pm(\hat{z}+\hat{x})$ and ±$\pm(\hat{y}+\hat{z})$ bonds respectively, as shown in Fig. 1(b-d).
@@
Quantum loop model: Hamiltonian— We now define a minimal Hamiltonian à la Rokhsar-Kivelson[14]. We first define potential energy with two contributions:
@@
These energy contributions encode proximity of bonds. For example, electrons come closer when bonds form an acute angle rather than an obtuse angle. This can impose an energy cost due to Coulomb repulsion, corresponding to a positive value of $V$. In a triangle loop, we have three bonds that are very close to one another. This could give rise to a local distortion with the three vertices coming closer to one another, with an associated energy cost $V^{\prime}$. We treat $V$ and $V^{\prime}$ as independent parameters that can each take positive or negative values.
@@
This Hamiltonian differs from previous studies that have adapted Rokhsar-Kivelson-like terms to a loop setting[5, 6]. There is a difference even at the level of the Hilbert space, due to the bending constraint at each site that arises from orbital character. This further constrains the kinetic energy term – a flipping process is only allowed if the initial and final configurations satisfy the bending constraint. The second key difference is in the form of the potential energy terms. Previous studies have defined potential energy in direct analogy with the original Rokhsar-Kivelson model, associating an energy cost with each flippable plaquette. Here, we have defined potential energy on physical grounds motivated by the geometry and physics of TMDs.
Numerical approach—  To solve for the ground state of Eq. 10, we first enumerate all loop coverings to construct the Hilbert space. This is a challenging task, with no known analytic solution. We use a numerical approach, working with finite clusters and periodic boundaries. We use a recursive branching algorithm to find all loop coverings that satisfy the bending constraint. The number of configurations grows rapidly with system size: 42 for 3x3, 250 964 for 5x4, 2 720 400 for 6x4. For concreteness, we present results for the 6x4 cluster below. We expect the same qualitative behaviour with larger sizes.
We construct the Hamiltonian as a 2 720 400 $\times$ 2 720 400 matrix, with flipping processes entering in off-diagonal terms. The resulting matrix is sparse, with $\sim 10^{-4}\%$ of the entries being non-zero. We use a Krylov-space based approach to find the lowest-lying eigenstates.
@@
Phase diagram—  We consider the Hamiltonian in Eq. 10 as a function of three parameters, $t$, $V$ and $V^{\prime}$. Using exact diagonalization, we find the ground state as a linear superposition of all loop covers. To determine the nature of the ground state, we examine (a) the loop cover(s) with the largest weight, and (b) dimer-dimer correlations in the ground state. We identify three phases:
@@
We identify the acute stripe phase with the 1T phase[15, 16, 17] seen in many materials. Bond correlations in the QLM ground state show a clear stripe-like pattern. This pattern is very similar to the variation of bond lengths seen in 1T materials – see Supplemental Material[18]. In the QLM ground state, a stronger bond is indicated by a higher likelihood of dimer occupancy. In materials, a stronger bond is one with a shorter bond length.
(ii) Obtuse stripe phase: This phase appears at positive values of $V$ where obtuse angles are favoured. Indeed, the obtuse stripe configuration maximizes the number of obtuse angles with all other configurations having fewer obtuse angles. They also allow for a high degree of dynamics with many flippable plaquettes. The ground state is a linear superposition with dominant weight from an obtuse stripe configuration.
(iii) Trimerized phase: We find a single-configuration-state that contains an arrangement of triangle-loops. On the 6$\times$4 lattice, we find a trimerized phase with alternating ‘up’ and ‘down’ triangles as shown in Fig. 3(b). On other sizes, we find configurations with all-up or all-down – see Supplemental Material[18]. Trimerized phases do not allow for flipping processes within our model. Dynamics may appear in a more general model with longer-range rearrangements.
A trimerized phase is known to appear in AVX2, where A=Li, Na and X=O,S,Se. Previous studies have described trimerization as arising from orbital ordering[19, 20, 21]. Our QLM places the trimerized phase within the broader context of loop model phases.
Impurity textures—  Our QLM description is a real-space approach based on short-ranged valence bond formation. This can be contrasted with momentum-space-based studies, e.g., where distortions are viewed as exciton-like instabilities[22, 23, 24] or as Peierls transitions[25, 26, 27]. To build a case for a real-space approach, we present a testable prediction. We consider a single impurity where a d1 transition metal substitutes at the site of the d2 metal atom. For example, we may have a Nb atom at a Mo site in 1T-MoS2. A d1 atom can only form a single valence bond. As a result, it cannot allow a loop to pass through. Rather, it serves as a loop-termination point.
@@
As a further test, we may consider a d3, rather than d1, impurity. This creates a T-junction rather than a loop termination point. This leads to precisely the same distortion patterns as d1, as shown in the Supplemental Material[18].
@@
Discussion—  We have proposed a QLM description for 1T TMDs, where distortions are driven by short-ranged valence bond formation. Loops emerge from d2 character and the directionality of t2g overlaps, with each loop containing a sequence of valence bonds. This picture is justified in a Mott insulator driven by strong intra-orbital Hubbard repulsion. If Hund’s coupling is weak, we may identify each valence bond as a covalent bond or a spin-singlet wavefunction. If Hund’s coupling is strong, two electrons on the same atom cannot form independent singlet bonds. Rather, electron spins will align to give rise to an effective spin-1 moment on the atom. Nevertheless, the QLM description can hold – if each loop is interpreted as an emergent Haldane chain. This picture has been evoked to describe the trimerized phase of LiVO2[20]. Based on the same idea, loop models have been proposed on the square lattice in the context of LaFeAsO[28] and more recently on the honeycomb lattice[29].
In MoX2 and WX2 (X=S,Se), the 1T structure is known to have a small electronic gap[26, 30]. However, the 1T structure is metallic[15]. Naively, this goes against our premise of localized electrons. If these distortion-free materials are described by a QLM, they may arise from a putative loop-liquid state. This state may host mobile charged excitations, reminiscent of holons and doublons in the resonating valence bond (RVB) theory of high-temperature superconductivity. This is an exciting direction for future studies, motivating the search for a quantum loop liquid in analogy with quantum spin liquids. Notably, a loop liquid phase has been seen in quantum Monte Carlo simulations of a related loop model[6].
@@
Supplemental Material for “Quantum loops in the 1T transition metal dichalcogenides”
A. Knowles,1 G. Baskaran,2,3,4 and R. Ganesh1
@@
2The Institute of Mathematical Sciences, CIT Campus, Chennai 600 113, India
3Department of Physics, Indian Institute of Technology Madras, Chennai 600036, India
4Perimeter Institute for Theoretical Physics, Waterloo, ON N2L 2Y5, Canada
(Dated: February 11, 2025)
@@ III Trimerized phases
Figure S5: a) Dimer density in the acute stripe ground state of the QLM. The ground state is projected to only retain states where dimers are present at the two bonds at the bottom left. Bond colour and line width are proportional to the dimer probability. b) Bond lengths in the MoS2 using data from Ref. 11. Darker, thicker bonds have shorter bond lengths while lighter, thinner bonds are longer.